# Emphasizing Discriminative Features for Dataset Distillation in Complex Scenarios

### v1

@@ Abstract
Dataset distillation has demonstrated strong performance on simple datasets like CIFAR, MNIST, and TinyImageNet but struggles to achieve similar results in more complex scenarios. In this paper, we propose EDF (emphasizes the discriminative features), a dataset distillation method that enhances key discriminative regions in synthetic images using Grad-CAM activation maps. Our approach is inspired by a key observation: in simple datasets, high-activation areas typically occupy most of the image, whereas in complex scenarios, the size of these areas is much smaller. Unlike previous methods that treat all pixels equally when synthesizing images, EDF uses Grad-CAM activation maps to enhance high-activation areas. From a supervision perspective, we downplay supervision signals that have lower losses, as they contain common patterns. Additionally, to help the DD community better explore complex scenarios, we build the Complex Dataset Distillation (Comp-DD) benchmark by meticulously selecting sixteen subsets, eight easy and eight hard, from ImageNet-1K. In particular, EDF consistently outperforms SOTA results in complex scenarios, such as ImageNet-1K subsets. Hopefully, more researchers will be inspired and encouraged to improve the practicality and efficacy of DD. Our code and benchmark will be made public at NUS-HPC-AI-Lab/EDF.
@@
1 Introduction
Dataset Distillation (DD) has been making remarkable progress since it was first proposed by Wang et al. (2020). Currently, the mainstream of DD is matching-based methods (Zhao et al., 2021; Zhao & Bilen, 2021b; a; Cazenavette et al., 2022), which first extract patterns from the real dataset, then define different types of supervision to inject extracted patterns into the synthetic data. On several simple benchmarks, such as CIFAR (Krizhevsky, 2009) and TinyImageNet (Le & Yang, 2015), existing matching-based DD methods can achieve lossless performance (Guo et al., 2024; Li et al., 2024).
However, there is still a long way to go before DD can be practically used in real-world applications, i.e., images in complex scenarios are characterized by significant variations in object sizes and the presence of a large amount of class-irrelevant information. To show that current DD methods fail to achieve satisfying performance in complex scenarios, we apply trajectory matching (Guo et al., 2024) on a 10-class subset from ImageNet-1K extracted by selecting similar classes of CIFAR-10, called IN1K-CIFAR-10. As depicted in Figure 1(a), under three compressing ratios, DD’s performances [^1] We compare recovery ratios because datasets are different. on IN1K-CIFAR-10 are consistently worse than those on CIFAR-10.
To figure out the reason behind the above results, we take a closer look at the ImageNet-1K and CIFAR-10 from the data perspective. One key observation is that the percentage of discriminative features in the complex scenario, which can be visualized by Grad-CAM (Selvaraju et al., 2016), is much lower. From Figure 1(b), CIFAR-10 images are sticker-like, and activation maps have higher means and larger highly activated areas. By contrast, activation maps of the IN1K-CIFAR-10 subset exhibit much lower activation means and smaller highly activated areas. Previous methods (Cazenavette et al., 2022; Guo et al., 2024) treat all pixels of synthetic images equally. Therefore, when applying these methods to more complex scenarios, the large ratio of low-activation areas leads to non-discriminative features dominating the learning process, resulting in a performance drop.
@@ 1 Introduction
From the supervision perspective, taking trajectory matching as an example, we investigate the impact of trajectories with different loss values on synthetic images. Specifically, we compare the trajectory-matching performance between i) using trajectory parameters with low losses only and ii) using trajectory parameters with high losses only. The effect on a single image is shown in Figure 2(a). Low-loss supervision reduces the mean of Grad-CAM activation maps and shrinks the high-activation area (also shifted). By contrast, high-loss supervision increases the activation mean and expands the high-activation region.
Additionally, we visualize the inter-class feature distribution for a broader view. In Figure 2(b), we show the t-SNE of features of synthetic images distilled by only low-loss trajectories. As the distillation proceeds, synthetic image features of different classes continuously come closer, and the confusion among classes becomes more severe, which is likely caused by common patterns. The above two phenomenons confirm that low-loss supervision primarily reduces the representation of discriminative features and embeds more common patterns into synthetic images, harming DD’s performance.
@@
In summary, EDF consistently achieves state-of-the-art (SOTA) performance across various datasets, underscoring the effectiveness of emphasizing discriminative features. On several ImageNet-1K subsets, EDF achieves lossless performance. To the best of our knowledge, we are the first to achieve lossless performance on ImageNet-1K subsets. We build the Complex Dataset Distillation benchmark based on complexity, providing convenience for future research to continue improving DD’s performance in complex scenarios.
@@ 2 Method
Our approach, Emphasize Discriminative Features (EDF), enhances discriminative features in synthetic images during distillation. As shown in Figure 3, EDF first trains trajectories on real images $T$ and synthetic images $S$ and computes the trajectory matching loss. Then, Common Pattern Dropout filters out low-loss supervision signals, retaining high-loss ones for backpropagation. After obtaining gradients for the synthetic images, Discriminative Area Enhancement uses dynamically extracted Grad-CAM activation maps to rescale pixel gradients, focusing updates on discriminative regions.
@@ 2.1 Common Pattern Dropout
To generate expert and student trajectories, we first train agent models on real data for $E$ epochs, saving the resulting parameters as expert trajectories, denoted by $\{\theta_{t}\}_{0}^{E}$. At each distillation iteration, we randomly select an initial point $\theta_{t}$ and a target point $\theta_{t+M}$ from these expert trajectories. Similarly, student trajectories are produced by initializing an agent model at $\theta_{t}$ and training it on the synthetic dataset, yielding the parameters $\{\hat{\theta}_{t}\}_{0}^{N}$. The trajectory matching loss is then computed by comparing the final student parameters $\hat{\theta}_{t+N}$ with the expert’s target parameters $\theta_{t+M}$, normalized by the initial difference:
@@ 2.2 Discriminative Area Enhancement
Grad-CAM generates class-specific activation maps by leveraging the gradients that flow into the final convolutional layer, highlighting key areas relevant for predicting a target class. To compute these maps, we first train a convolutional model $\mathcal{G}$ on the real dataset. Following the Grad-CAM formulation (Equation 3), we calculate the activation map for each class $c$: $M^{c}\in\mathbb{R}^{IPC\times H\times W}$ on the synthetic images (IPC is the number of images per class). The activation map $M^{c}$ is a gradient-weighted sum of feature maps across all convolutional layers:
$\alpha^{c}=\frac{1}{Z}\sum_{h}\sum_{w}\frac{\partial y^{c}}{\partial A^{l}_{h,w}}\ \ \ \ \ \ \ \ \ \ M^{c}=ReLU(\sum_{l}\alpha^{c}_{l}A^{l}),$ (3)
$\alpha^{c}$ represents the weight of the activation of the $l$-th convolutional layer, $A^{l}$, computed by gradients. Finally, we concatenate $M^{c}$ of all images in class $c$ and obtain $M\in\mathbb{R}^{|S|\times H\times W}$.
@@
A major limitation of previous DD algorithms (Cazenavette et al., 2022; Du et al., 2022; Guo et al., 2024) on the complex scenario is that they treat each pixel equally and provide no guidance for the distillation process on which area of synthetic images should be emphasized. Therefore, we propose to update synthetic images in a biased manner. Instead of treating each pixel equally, we enhance the significance of discriminative areas by guiding the optimization with activation maps extracted in the previous step. We define the discriminative area of a synthetic image as the percentage of pixels with activation values above the mean since synthetic images are dynamically changing (see Section 4.3 for discussion). Specifically, we process activation maps from the previous step with a function $\mathcal{F}(M,\beta)$ to create weights for pixel gradients as follows:
@@
$(\nabla D_{syn})_{edf}=\nabla D_{syn}\circ\mathcal{F}(M,\beta).$ (5)
@@
As synthetic images are optimized, high-activation regions shift over time. To capture these changes, we recompute the activation maps every $K$ iterations, focusing updates on the most relevant areas. The frequency $K$ is a tunable hyperparameter, adjusted based on the learning rate of the synthetic images (see Section 4.3 for details). This ensures evolving discriminative areas are accurately captured. The complete algorithm is provided in the appendix A.
3 Complex DD Benchmark
@@
We introduce the Complex Dataset Distillation (Comp-DD) benchmark, which is constructed by selecting subsets from ImageNet-1K based on their complexity. This benchmark represents an early and pioneering effort to address dataset distillation in complex scenarios. Although there are numerous benchmarks (Krizhevsky, 2009; Le & Yang, 2015; Cui et al., 2022b) for simpler tasks, there is a notable absence of benchmarks designed specifically for complex scenarios. This gap presents a significant challenge to advancing research in this area and limits the practical application of dataset distillation. To bridge this gap, we propose the first dataset distillation benchmark explicitly built around scenario complexity, aiming to promote further exploration within the DD community.
@@
We evaluate the complexity of an image by measuring the average size of high-activation regions of the Grad-CAM activation map. Using a pre-trained ResNet model, we first generate Grad-CAM activation maps for all images, class by class. For each image, we calculate the percentage of pixels with activation values above a predefined threshold (set to 0.5 in our case), with higher percentages indicating lower complexity (more clarifications can be found in Appendix D.2). Formally, the complexity of the $i$-th image is computed as $1-\frac{\sum_{h}\sum_{w}\mathbbm{1}[M^{i}_{h,w}\geq 0.5]}{H\cdot W}$ where $\mathbbm{1}$ is the indicator function. The complexity of each class is then determined by averaging the complexity scores across all images within that class.
@@
To reduce the influence of class differences, we select subsets from each category, where a category consists of classes representing visually similar objects or animals of the same species. This approach allows us to focus on complexity while controlling for inter-class variability.
Specifically, we first manually identify representative categories in ImageNet-1K with sufficient numbers of classes ($\geq 20$). For each category, we rank the classes by complexity in descending order. Following established practice, we construct two ten-class subsets for each category: the easy subset, comprising the ten least complex classes, and the hard subset, comprising the ten most complex classes. The subset-level complexity is determined by averaging the complexity scores across all classes in each subset.
@@
We carefully selected eight categories from ImageNet-1K: Bird, Car, Dog, Fish, Snake, Insect, Round, and Music. Each category contains two ten-class subsets: one easy and one hard, with difficulty determined by the complexity metrics outlined above. Figure 4(a) summarizes the number of training images in each subset, while all subsets contain 500 images in the validation set. To illustrate the difference between easy and hard subsets, Figure 4(b) compares the subset-level complexity for each category. As expected, the hard subsets display significantly higher complexity than the easy subsets. For a detailed breakdown of the classes in each subset, please refer to Appendix D.1.
4 Experiment
@@ 4.1 Experimental Setup
We conduct a comprehensive evaluation of EDF on six ten-class subsets (Howard, 2019) of ImageNet-1K (ImageNette, ImageWoof, ImageMeow, ImageYellow, ImageFruit, and ImageSquawk) and a one-hundred-class subset (ImageNet100). Each subset contains ten classes, with approximately 13,000 images in the training set and 500 images in the validation set. On the Comp-DD benchmark, we report the results of the Bird, Car, and Dog categories. For all experiments, we use a 5-layer ConvNet (ConvNetD5) as both the distillation and the evaluation architecture. For cross-architecture evaluation (see results in Appendix C.1), we validate synthetic data accuracy on Alexnet (Krizhevsky et al., 2012), VGG11 (Simonyan & Zisserman, 2014), and ResNet18 (He et al., 2015).
@@
We compare two baselines: dataset distillation (DD) methods and methods utilizing knowledge distillation (Eval. w/ Knowledge Distillation) (Hinton et al., 2015). For DD methods, we include trajectory-matching-based methods such as MTT (Cazenavette et al., 2022), FTD (Du et al., 2022), and DATM (Guo et al., 2024). In the knowledge distillation group, we compare against SRe2L (Yin et al., 2023) and RDED (Sun et al., 2023). The results for subsets not covered in these papers are obtained through replication using the official open-source codebases and hyperparameters.
@@ 4.2 Main Results
ImageNet-1K Subsets.
We mainly conduct extensive experiments on various ImageNet-1K subsets and compare the performance of EDF with other different approaches. The detailed results are shown in Table 1. EDF consistently achieves state-of-the-art (SOTA) results across all settings when compared to other dataset distillation methods. On larger IPCs, i.e., 200 or 300, the performance of EDF significantly outperforms that observed with smaller IPCs.
Table 2: Bolded entries are lossless performances under IPC300.
@@
We achieve lossless performances on ImageMeow and ImageYellow under IPC300, 23% of real data, as shown in Table 2. When evaluated against Eval. w/ Knowledge Distillation methods, our distilled datasets outperform SRe2L and RDED in 14 out of 18 settings. It is important to note that applying knowledge distillation (KD) for evaluation tends to reduce EDF’s pure dataset distillation performance, particularly in low IPC (images per class) settings such as IPC1 and IPC10. This occurs because smaller IPCs lack the capacity to effectively incorporate the knowledge from a well-trained teacher model. We also provide results without knowledge distillation in Appendix C.2.
@@
The results for EDF on the Bird, Car, and Dog categories from the Comp-DD Benchmark are shown in Table 3. EDF demonstrates superior test accuracies and recovery ratios across both easy and hard subsets. As expected, the recovery ratios for easy subsets are consistently higher than those for hard subsets, confirming that the hard subsets present a greater challenge for dataset distillation methods. These results validate our complexity metrics, which effectively distinguish the varying levels of difficulty between easy and hard subsets.

### v2

@@ Abstract
Dataset distillation has demonstrated strong performance on simple datasets like CIFAR, MNIST, and TinyImageNet but struggles to achieve similar results in more complex scenarios. In this paper, we propose EDF (emphasizes the discriminative features), a dataset distillation method that enhances key discriminative regions in synthetic images using Grad-CAM activation maps. Our approach is inspired by a key observation: in simple datasets, high-activation areas typically occupy most of the image, whereas in complex scenarios, the size of these areas is much smaller. Unlike previous methods that treat all pixels equally when synthesizing images, EDF uses Grad-CAM activation maps to enhance high-activation areas. From a supervision perspective, we downplay supervision signals produced by lower trajectory-matching losses, as they contain common patterns. Additionally, to help the DD community better explore complex scenarios, we build the Complex Dataset Distillation (Comp-DD) benchmark by meticulously selecting sixteen subsets, eight easy and eight hard, from ImageNet-1K. In particular, EDF consistently outperforms SOTA results in complex scenarios, such as ImageNet-1K subsets. Hopefully, more researchers will be inspired and encouraged to improve the practicality and efficacy of DD. Our code and benchmark have been made public at NUS-HPC-AI-Lab/EDF.
00footnotetext: equal contribution (kai.wang@comp.nus.edu.sg, lizekai@u.nus.edu)  advising contribution. Part of this work was completed at CMU.
1 Introduction
Dataset Distillation (DD) has been making remarkable progress since it was first proposed by [54]. Currently, the mainstream of DD is matching-based methods [71, 69, 68, 1], which first extract patterns from the real dataset, then define different types of supervision to inject extracted patterns into the synthetic data. On several simple benchmarks, such as CIFAR [19] and TinyImageNet [21], existing DD methods can achieve lossless performance [13, 25].
@@ 1 Introduction
However, DD still struggles to be practically used in real-world applications, i.e., images in complex scenarios are characterized by significant variations in object sizes and the presence of a large amount of class-irrelevant information. To show that current DD methods fail to achieve satisfying performance in complex scenarios, we apply trajectory matching [13] on a 10-class subset from ImageNet-1K extracted by selecting similar classes of CIFAR-10, called IN1K-CIFAR-10. As depicted in Figure 1(a), under three compressing ratios, DD’s performances [^1] We compare recovery ratios because datasets are different. on IN1K-CIFAR-10 are consistently worse than those on CIFAR-10.
To figure out the reason, we take a closer look at the ImageNet-1K and CIFAR-10 from the data perspective. One key observation is that the percentage of discriminative features in the complex scenario, which can be visualized by Grad-CAM [42], is much lower. From Figure 1(b), CIFAR-10 images are sticker-like, and activation maps have higher means and larger highly activated areas. By contrast, activation maps of the IN1K-CIFAR-10 subset exhibit much lower activation means and smaller highly activated areas. Previous methods [1, 13] treat all pixels of synthetic images equally. Therefore, when applying these methods to more complex scenarios, the large ratio of low-activation areas leads to non-discriminative features dominating the learning process, resulting in a performance drop.
@@
As for supervision, taking trajectory matching as an example, we investigate the impact on synthetic images of supervision signals brought by different trajectory-matching losses. Specifically, we compare the performance between i) using only supervision from low trajectory-matching losses to update synthetic images and ii) using only supervision from high trajectory-matching losses to update synthetic images. The effect on a single image is shown in Figure 2(a). Low-loss supervision reduces the mean of Grad-CAM activation maps and shrinks the high-activation area (also shifted). By contrast, high-loss supervision increases the activation mean and expands the high-activation region.
Additionally, we visualize the inter-class feature distribution for a broader view. In Figure 2(b), we show the t-SNE of features of synthetic images distilled by only low trajectory-matching losses. As the distillation proceeds, synthetic image features of different classes continuously come closer, and the confusion among classes becomes more severe, which is likely caused by common patterns. The above two phenomenons confirm that low-loss supervision primarily reduces the representation of discriminative features and embeds more common patterns into synthetic images, harming DD’s performance.
@@
In summary, EDF consistently achieves state-of-the-art (SOTA) performance across various datasets, underscoring its effectiveness. On some ImageNet-1K subsets, EDF achieves lossless performance. To the best of our knowledge, we are the first to achieve lossless performance on ImageNet-1K subsets. We build the Complex Dataset Distillation benchmark based on complexity, providing convenience for future research to continue improving DD’s performance in complex scenarios.
@@ 2 Method
Our approach, Emphasize Discriminative Features (EDF), enhances discriminative features of synthetic images. As shown in Figure 3, EDF first trains trajectories on real images $T$ and synthetic images $S$ and computes the trajectory matching loss. Then, Common Pattern Dropout filters out low-loss supervision signals, retaining high-loss ones for backpropagation. After obtaining gradients for the synthetic images, Discriminative Area Enhancement uses dynamically extracted Grad-CAM activation maps to rescale pixel gradients, focusing updates on discriminative regions. Although EDF is built on the trajectory-matching backbone, the two modules can be easily applied to other matching-based methods with minor modifications.
@@ 2.1 Common Pattern Dropout
To generate expert and student trajectories, we first train agent models on real data for $E$ epochs, saving the resulting parameters as expert trajectories, denoted by $\{\theta_{t}\}_{0}^{E}$. At each distillation iteration, we randomly select an initial point $\theta_{t}$ and a target point $\theta_{t+M}$ from these expert trajectories. Similarly, student trajectories are produced by initializing an agent model at $\theta_{t}$ and training it on the synthetic dataset, yielding the parameters $\{\hat{\theta}_{t}\}_{0}^{N}$. The trajectory matching loss is the distance between final student parameters $\hat{\theta}_{t+N}$ and expert’s target parameters $\theta_{t+M}$, normalized by the initial difference:
@@ 2.2 Discriminative Area Enhancement
Grad-CAM generates class-specific activation maps by leveraging the gradients that flow into the final convolutional layer, highlighting key areas relevant to predicting a target class. To compute these maps, we first train a convolutional model $\mathcal{G}$ on the real dataset. Following the Grad-CAM formulation (Equation 3), we calculate the activation map for each class $c$: $M^{c}\in\mathbb{R}^{IPC\times H\times W}$ on the synthetic images (IPC is the number of images per class). The activation map $M^{c}$ is a gradient-weighted sum of feature maps across all convolutional layers:
$\alpha^{c}=\frac{1}{Z}\sum_{h}\sum_{w}\frac{\partial y^{c}}{\partial A^{l}_{h,w}},\quad M^{c}=ReLU(\sum_{l}\alpha^{c}_{l}A^{l}).$ (3)
$\alpha^{c}$ represents the weight of activation corresponding to the $l$-th feature map, $A^{l}$, computed by gradients. Finally, we concatenate $M^{c}$ of all images in class $c$ and obtain $M\in\mathbb{R}^{|S|\times H\times W}$. Note that the capability of the model used to extract activation maps does not affect the performance. We provide ablation in the Appendix 11.5.
@@
A major limitation of previous DD algorithms [1, 8, 13] on the complex scenario is that they treat each pixel equally and provide no guidance for the distillation process on which area of synthetic images should be emphasized. Therefore, we propose to update synthetic images in a biased manner. Instead of treating each pixel equally, we enhance the significance of discriminative areas by guiding the optimization with activation maps extracted in the previous step. We define the discriminative area of a synthetic image as the percentage of pixels with activation values above the mean since synthetic images are dynamically changing (see Section 4.3 for discussion). Specifically, we process activation maps from the previous step with a function $\mathcal{F}(M,\beta)$ to create weights for pixel gradients as follows:
@@
$(\nabla D_{syn})_{edf}=\nabla D_{syn}\odot\mathcal{F}(M,\beta).$ (5)
@@
As synthetic images are optimized, high-activation regions shift over time. To capture these changes, we recompute the activation maps every $K$ iterations, focusing updates on the most relevant areas. The frequency $K$ is a tunable hyperparameter, adjusted based on the learning rate of the synthetic images (see Section 4.3 for details). This ensures evolving discriminative areas are accurately captured. The complete algorithm is provided in the Appendix 9.
3 Complex Dataset Distillation Benchmark
@@
We introduce the Complex Dataset Distillation (Comp-DD) benchmark, which is constructed by selecting subsets from ImageNet-1K based on their complexity. This benchmark represents an early and pioneering effort to address dataset distillation in complex scenarios. Although there are numerous benchmarks [19, 21, 7] for simpler tasks, there is a notable absence of benchmarks designed specifically for complex scenarios. This gap presents a significant challenge to advancing research in this area and limits the practical application of dataset distillation. To bridge this gap, we propose the first dataset distillation benchmark explicitly built around scenario complexity, aiming to promote further exploration within the DD community.
@@
We evaluate the complexity of an image by measuring the average size of high-activation regions of the Grad-CAM activation map. Using a pre-trained ResNet model, we first generate Grad-CAM activation maps for all images, class by class. For each image, we calculate the percentage of pixels with activation values above a predefined threshold $\alpha$ (set to 0.5 in our case), with higher percentages indicating lower complexity (more clarifications can be found in Appendix 12.2). Formally, the complexity of the $i$-th image is computed as $1-\frac{\sum_{h}\sum_{w}\mathbbm{1}[M^{i}_{h,w}\geq\alpha]}{H\cdot W}$ where $\mathbbm{1}$ is the indicator function. The complexity of each class is then determined by averaging the complexity scores across all images within that class.
@@
To reduce the influence of class differences, we select subsets from each category, where a category consists of classes representing visually similar objects or animals of the same species. This approach allows us to focus on complexity while controlling for inter-class variability. Specifically, we first manually identify representative categories in ImageNet-1K with sufficient numbers of classes ($\geq 20$). For each category, we rank the classes by complexity in descending order. Following established practice, we construct two ten-class subsets for each category: the easy subset, comprising the ten least complex classes, and the hard subset, comprising the ten most complex classes. The subset-level complexity is determined by averaging the complexity scores across all classes in each subset.
@@
We carefully selected eight categories from ImageNet-1K: Bird, Car, Dog, Fish, Snake, Insect, Round, and Music. Each category contains two ten-class subsets: one easy and one hard, with difficulty determined by the complexity metrics outlined above. Figure 4(a) summarizes the number of training images in each subset, while all subsets contain 500 images in the validation set. To illustrate the difference between easy and hard subsets, Figure 4(b) compares the subset-level complexity for each category. As expected, the hard subsets display significantly higher complexity than the easy subsets. For a detailed breakdown of the classes in each subset, please refer to Appendix 12.1.
4 Experiments
@@ 4.1 Experimental Setup
We conduct a comprehensive evaluation of EDF on six ten-class subsets [16] of ImageNet-1K (ImageNette, ImageWoof, ImageMeow, ImageYellow, ImageFruit, and ImageSquawk) and a one-hundred-class subset (ImageNet100). Each subset contains ten classes, with approximately 13,000 images in the training set and 500 images in the validation set. On the Comp-DD benchmark, we report the results of the Bird, Car, and Dog categories. For all experiments, we use a 5-layer ConvNet (ConvNetD5) as both the distillation and the evaluation architecture. For cross-architecture evaluation (see results in Appendix 11.1), we validate synthetic data accuracy on Alexnet [20], VGG11 [45], ResNet18 [14].
@@
We compare two baselines: dataset distillation (DD) methods and methods utilizing knowledge distillation (Eval. w/ Knowledge Distillation) [15]. For DD methods, we compare trajectory-matching-based methods such as MTT [1], FTD [8], and DATM [13]. In the knowledge distillation group, we compare against SRe2L [64] and RDED [48]. The results for subsets not covered in these papers are obtained through replication using the official open-source recipe.
@@ 4.2 Main Results
Table 2: Bolded entries are lossless performances under IPC300.
Dataset IPC DD Eval. w/ KD
DATM EDF RDED EDF
CIFAR-10 1 46.9±0.5 47.8±0.3 23.5±0.3 28.1±0.5
50 76.1±0.3 77.3±0.5 68.4±0.1 69.5±0.3
500 83.5±0.2 84.8±0.5 / 86.1±0.4
Tiny-ImageNet 1 17.1±0.3 17.8±0.2 12.0±0.1 10.3±0.4
10 31.1±0.3 32.5±0.5 39.6±0.1 40.5±0.7
50 39.7±0.3 41.1±0.6 47.6±0.2 48.9±0.3
Table 3: EDF also outperforms baselines of two approaches on datasets in simple scenarios, such as CIFAR-10 and Tiny-ImageNet.
ImageNet-1K Subsets.
We compare our approach with previous works on various ImageNet-1K subsets. As shown in Table 1, EDF consistently achieves state-of-the-art (SOTA) results across all settings. On larger IPCs, i.e., 200 or 300, the performance of EDF significantly outperforms that observed with smaller IPCs. We achieve lossless performances on ImageMeow and ImageYellow under IPC300, 23% of real data, as shown in Table 2. When evaluated against Eval. w/ Knowledge Distillation methods, our distilled datasets outperform SRe2L and RDED in most settings. It is important to note that applying knowledge distillation (KD) for evaluation tends to reduce EDF’s pure dataset distillation performance, particularly in low IPC (images per class) settings such as IPC1 and IPC10. This occurs because smaller IPCs lack the capacity to effectively incorporate the knowledge from a well-trained teacher model. We also provide results without knowledge distillation in Appendix 11.2.
CIFAR-10 and Tiny-ImageNet.
We also evaluate our approach on simple datasets, i.e., CIFAR-10 and Tiny-ImageNet. As illustrated in Table 3, EDF can still outperform the SOTAs of the DD and Eval. w/ KD approaches in most settings. The results confirm the good generalization of our approach in simple scenarios.
@@
The results for EDF on the Bird, Car, and Dog categories from the Comp-DD Benchmark are shown in Table 4. EDF demonstrates superior test accuracy. As expected, the recovery ratios for easy subsets are consistently higher than those for hard subsets, confirming that the hard subsets present a greater challenge for DD methods. These results validate our complexity metrics, which effectively distinguish the varying levels of difficulty between easy and hard subsets.